OpenStudy (anonymous):

Explain, in complete sentences, how you would use the two methods (distribution and box method) to find the product of (x – 2) (x – 3). Please include your steps for each method and the resulting product. Which method do you prefer and why? I need help with the box method part. Idk how to do it on the computer?? Can someone help me with this?? Please and thank you :)

OpenStudy (anonymous):

This is what I have so far: I prefer the distribution method because, first of all, I find it easier, and secondly, that is how Mrs. Huerkamp explained it in the live lesson. To do this you simply utilize the FOIL technique by multiplying the first, outer, inner and last variables together. (x) (x) = x2 (x) (-3) = -3x (-2) (x) = -2x (-2) (-3) = 6 so your answer is: x2 - 3x - 2x + 6 = x2 - 5x + 6 The box method would be you make a 2x2 box and go up, down, and sideways to multiply the variables. It’s the same basic principal, just a little different:

hero (hero):

Well...

hero (hero):

It would be easier to say that for distribution method, recall that the formula for distributive property is a(b+c) = ab + bc. In this case: a = (x-2) b = x c = -3 Therefore: a(b+c) = ab + bc (x-2)(x-3) = (x-2)x + (x-2)-3 Thus: (x-2)(x-3) = (x-2)x + (x-2)-3 = x^2 - 2x - 3x + 6 = x^2-5x + 6

OpenStudy (anonymous):

Ohhh, okay. Sorey for the late reply, I keep losing connection but what about the box method? How would I wrote that out on a computer? I can't get it to look right...

OpenStudy (anonymous):

sorry*

OpenStudy (anonymous):

write*

OpenStudy (anonymous):

I don't have the option of making a chart and i'm using a word document... and it has to be done online but i'm not sure how to explain it and draw it out on the computer?

hero (hero):

If you're using a word document, creating the box method is relatively easy, assuming you have MS Word 2010 or higher.